# Three-Dimensional to Layered Halide Perovskites: A Parameter-Free Hybrid Functional Method for Predicting Electronic Band Gaps

### v1

@@ Abstract
Accurately predicting electronic band gaps in halide perovskites using ab initio density functional theory (DFT) is essential for their application in optoelectronic devices. Standard hybrid functionals such as HSE and PBE0 can overcome the limitations of DFT with reasonable computational cost but are known to underestimate the measured band gaps for layered halide perovskites. In this study, we assess the performance of the doubly screened dielectric-dependent hybrid (DSH) functional for predicting band gaps in three-dimensional (3D) and layered hybrid perovskites. We show that the DSH functional, which employs material-dependent mixing parameters derived from macroscopic dielectric constants, provides accurate band gap predictions for 3D halide perovskites when structural local disorder is considered. For layered hybrid perovskites, we identified a challenge in setting the relevant screening parameter, and we found that the DSH functional based on average dielectric constants overestimates the band gaps. To improve the predictions and stay in a parameter-free ab initio workflow, we propose to use the calculated dielectric constant of the respective 3D perovskites. We find that the PBE0 and DSH functionals using the dielectric constants of the 3D perovskite accurately predict experimental gaps, with mean absolute errors of 0.21 eV and 0.14 eV, respectively, for layered perovskites with various organic spacers, as well as for multilayered with n = 2, 3. Notably, the HSE functional systematically underestimates the band gaps in layered perovskites, with absolute errors of 0.84 eV. We attribute the root of this failure to the absence of non-local long-range dielectric screening, a critical factor for halide perovskites. The computational framework introduced here provides an efficient parameter-free ab initio methodology for predicting the electronic properties of 3D and layered halide perovskites and their heterostructures, aiding in developing advanced optoelectronic devices.
@@
Lead-based hybrid halide perovskites have emerged as versatile materials with tunable structural, electronic, and optical properties due to their unique inorganic corner-sharing octahedral lattices and complex organic components. These diverse material compositions enable a spectrum of functionalities that have been successfully applied in various technological fields, such as solar cells, photodetectors, and light-emitting devices 1, 2, 3, 4. Halide perovskites are generally classified into three-dimensional (3D) and layered halide perovskites. The 3D halide perovskites are represented by the general formula AMX3, where M is typically a divalent cation in the center of the octahedron such as Pb2+ or Sn2+, X is a halogen ( I, Br, or Cl), and the A site is typically occupied by Cs+ or organic cations such as methylammonium (MA) or formamidinium (FA). In recent years, solar cells based on 3D halide perovskites have achieved remarkable power conversion efficiencies today comparable to those achieved by silicon cells 1. However, these 3D materials can suffer from poor environmental stability when exposed to air and/or illumination 5. Their layered counterparts, also called two-dimensional (2D) halide perovskites, are further classified into Ruddlesden-Popper (RP), Dion-Jacobson (DJ), and alternating cation (ACI) structures, depending on the choice of organic molecules that separate the layers and the resulting stacking arrangement 6. For example, the RP structure typically has two monovalent organic cations and the general formula , where is the number of connected MX6 octahedra in the structure 7. Layered halide perovskites can exhibit significantly higher material stability than 3D perovskites. However, their photovoltaic performance is usually lower 8, 9. Within these layered materials, the organic moieties are usually insulating; thus, the charge carriers are primarily constrained to the inorganic layers. This leads to the natural formation of quantum wells (QW), in which the optical band gap increases with decreasing due to quantum and dielectric confinement 10, 11. Most recently, the combination of the two in the so-called 3D/2D heterostructures has achieved superior stability compared to 3D and higher efficiency than layered halide perovskites, making them a promising platform for optimizing device performance in various architectures 12. Consequently, knowledge about the energy levels is critical to monitor the band alignment between layered and 3D structures 12.
@@
Density Functional Theory (DFT) calculations, employing semi-local exchange-correlation (XC) functionals like the generalized gradient approximation (GGA) with spin-orbit coupling effects, lead to a sizable and systematic underestimation of the electronic band gap. The state-of-the-art method to go beyond DFT is Hedin’s many-body quasiparticle perturbation theory 13 that accounts for exchange-correlation effects via the self-energy, defined based on the Green’s function, , and the dynamical Coulomb screening potential . The quasiparticle conduction band minimum (CBM) and valence band maximum (VBM) can be interpreted as the ionization energy (IE) measured by ultraviolet photoemission spectroscopy (UPS) and the electron affinity (EA) obtained from inverse photoemission spectroscopy (IPES), respectively. However, the approximation is computationally expensive 14, which limits its use in large systems containing hundreds or thousands of electrons, such as layered halide perovskites and very few works have reported the GW band gaps in the entire structure (i.e., including the organic moieties) of layered halide perovskites 15, 16, 17, 18, 19.
@@
An alternative approach to is using the so-called hybrid functionals within DFT constructed by admixing a non-local Hartree fraction to the local exchange potential 20. This mixing mitigates the self-interaction error, and has been shown for several semiconductors to result in improved band gaps and reasonable energy level alignment at a computational cost that can be significantly smaller than calculations 21. Well-known hybrid functionals include PBE0 22, which adds 25% (i.e., the mixing parameter is set to 0.25) exact Hartree-Fock exchange to the exchange-correlation potential. HSE 23, 24 is similar to PBE0 but introduces a range-separation parameter that partition the Coulomb interaction into short- and long-range components. These hybrid functionals significantly improve standard DFT, yet PBE0 tends to overestimate band gaps, particularly for narrow band gap materials. At the same time, HSE underestimates band gaps of moderate band gap semiconductors and wide band insulators 25. To overcome this limitation, the mixing parameters and are often modified to reproduce experimental results, which limits the predictive capability of hybrid functionals 26, 27, 28. A good example of this limitation is the case of halide perovskites, for which hybrid functionals are being employed with a wide range of mixing parameters, in many cases surpassing the 60% of exact exchange 27, 29, 28, 30. Such manual tuning of the parameters in hybrid functionals is impractical and ineffective if the band gap of the material is not known a priori. It can also be challenging to predict suitable mixing parameters in complex heterostructures. Hence, it is desirable to develop mixing parameters independent of empirical tuning and have an intrinsic physical meaning.
@@
To this end, it has been shown that the inverse macroscopic dielectric constant , which relates to the screening strength in materials, can be employed in a "parameter-free" dielectric-dependent hybrid functional (DDH) which gives good agreement with experiments for representative materials with a wide range of band gaps and dielectric constants 31, 32, 33. A fully self-consistent version sc-DDH, which involves evaluating at the end of each DDH cycle, was also developed 34. A range-separated version (sc-RS-DDH) 35, with the long-range screening set at the empirical value of 0.25, has been proposed that further improves the band gap prediction with respect to sc-DDH. In the same spirit, but without any empirical parameter, two types of dielectric-dependent hybrid functionals have recently been introduced. First, the doubly screened dielectric-dependent hybrid (DSH) functional 36, which accounts for both electronic and metallic screening, and the dielectric-dependent range-separated hybrid functional using the Coulomb attenuated method (DD-RS-CAM) 37. Although both functionals use the same model dielectric function 38, 39, in each case the range separation parameter is determined slightly differently. In DSH, is derived from modified Thomas-Fermi screening, whereas in DD-RS-CAM is obtained by fitting the first-principles RPA dielectric function with the model dielectric function. Nevertheless, these two functionals are practically the same and do not give rise to significant differences in the predicted band gaps of materials 35, 40. The model macroscopic dielectric function in DSH and DD-RS-CAM disregards both local field effects (i.e., off-diagonal elements) and dynamical effects and can be viewed as a static approximation to the diagonal Coulomb hole and screened exchange (COHSEX) 13, 41 approximation to the GW self-energy. Overall, the validity of this approximation depends on how well the semi-local correlation represents the Coulomb hole and the significance of the dynamical effects. 37 Table S1 summarizes the parameters of the different hybrid functionals.
@@
To date, DSH has been applied to 3D halide perovskites 42, 43, 40, 44, and, to the best of our knowledge, no reports are validating DSH for the prediction of the band gap of layered halide perovskites, which remains a challenge. Nevertheless, the computational studies on 3D perovskites employ high-symmetry structures and disregard effects arising from structural local disorder, which are critical for the description of halide perovskites, playing an important role on their structural, electronic, optical, and phonon properties 45, 46, 47. In this work, we compute the electronic band gaps of the most prominent 3D and layered halide perovskites using the DSH hybrid functional. We do so to explore the possibility of developing a consistent workflow that would permit calculating the band gaps of all these materials without having to tune any parameter of the hybrid functionals empirically. To this end, we consider a set of layered halide perovskites, known to pose a challenge for hybrid functionals, containing different organic cation spacers and different halogen atoms (i.e., Cl, Br, and I), as well as a layered perovskite series with n = 2 and n=3. All materials studied in this work are shown in Figure 1. We assess the performance of the DSH functional in both its one-shot (DSH0) and self-consistent (sc-DSH) forms, comparing the results with those obtained using typical hybrid functionals, such as HSE and PBE0, and experimental measurements. We also explore the importance of the hybrid functionals’ short- and long-range separation on band gap for each system.
@@
We begin by analyzing how the doubly screened hybrid (DSH) functional performs on the prediction of the band gaps of prototypical cubic structures of 3D halide perovskites CsPbX3 (X = Cl, Br, I). We model the structures using the lattice parameters that are summarized in Table S2. All computed band gaps include spin-orbit coupling effects, important in halide perovskites 48. We compare these predictions with those obtained using PBE, PBE0, and HSE06, as well as the measured experimental electronic band gaps (See Table S2). Figure 2 shows the band gaps calculated employing different functionals (bars) and the measured band gaps (solid lines). As an additional reference, we include the standard semiconductors Si and GaAs, which have similar band gaps yet dielectric constants much higher than CsPbX3. Not surprisingly, all hybrid functionals drastically improve the PBE band gap values. For the standard semiconductors, HSE values are very close to the experimental band gaps, and so are the values obtained from the dielectric-dependent DSH0 and DSH. On the other hand, PBE0 significantly overestimates the band gap as range separation is critical for the functional form in both Si and GaAs. Moving on to the cubic phases of the 3D halide perovskites CsPbX3, the hybrid functionals fail to describe the band gaps as all functionals significantly underestimate the measured band gaps. However, this underestimation is related to neglecting local disorder (also referred to as positional "polymorphism"), which is critical for the accurate description of the electronic structure of halide perovskites at higher temperatures, including their dielectric properties 49, 45, 46, 47. Some of us have previously shown that such local disorder (shown in the inset structures of Figure 2) has significant effects on the band gaps, lattice dynamics, and electron-phonon coupling of halide perovskites 46, 47. Therefore, we consider the locally disordered configurations for CsPbX3 (p-CsPbX3) to assess the performance of the hybrid functionals. As shown in the last panel of Figure 2, we find a significant increase of the band gap values resulting from the local disorder for all the functionals, ranging from 570 meV (PBE, p-CsPbCl3) to 960 meV (DSH0, CsPbI3), leading to a better agreement with the measurements.
@@
Next, we discuss the effects of self-consistency on improving DSH band gaps. In both full-range 34 and range-separated dielectric dependent hybrid functionals 36, 37, early results indicated that self-consistency improves the band gaps values compared to one-shot methods, especially for narrow band gap materials 36. However, such self-consistency worsens the results for transition metal oxides, which is attributed to the underestimated dielectric constants 50. Here, we observe that self-consistency leads to relatively minor changes in the calculated band gaps, with the change being marginal in the case of Si, GaAs, and p-CsPbI3. Figure S3 shows the band gap variation and dielectric constants for each DSH iteration. We also note that the method employed to evaluate in the self-consistency cycle can be critical. can be computed using different methods, including density functional perturbation theory (DFPT) 51, finite electric field in the Berry phase approximation 52, or linear optics in the Random Phase Approximation (RPA) 53, each with different advantages and limitations discussed in the supporting information file. Here, we employ the finite electric field method, as detailed in the computational details, to include local field effects, which are important for halide perovskites. However, converging remains challenging when employing hybrid functionals.
@@
Next, we employ the DSH functional to set up a parameter-free workflow for layered halide perovskite materials. To this end, we consider a series of prototypical lead-based layered structures with various halogen atoms (i.e., Cl, Br, I) and different organic spacers, as seen in Table S3. Figure 1d-g shows the materials with the organic cations BA (butylammonium), PMA (phenylmethyl ammonium), PEA (phenylethylammonium) and FPEA (4F-phenylethylammonium), respectively. We have not considered locally disordered structures for the layered perovskites, as our calculations are based on experimental low-symmetry structures with significant distortion and therefore, the effects are expected to be small. For example, the change in total energy due to structural local disorder for orthorhombic CsPbI3 is reported at just 1 meV per formula unit 46. Also, in the absence of experimental electronic band gaps, we correct the experimental optical band gaps (E) using the empirical relation of the exciton binding energy (), derived for layered perovskites (i.e., ). 54 The ab initio computation of the exciton binding energy would require to solve the GW plus Bethe-Salpeter equation (GW+BSE). To date, only a handful of studies have used the GW+BSE method to compute band gap and exciton binding energy in the full structure (i.e., including organic moieties) of layered halide perovskites 15, 16, 17, 18, 19. For example, Yin et al. 15 investigated the origin of light emission in two layered perovskites. Giorgi et al. 16 calculated the optoelectronic properties of isolated and layered with n=1 and n=2. Filip et al. 18 analyzed the effects of various organic cations on the band gaps, dielectric screening, and exciton binding energies of a set of A2PbBr4 compounds.
@@
For the DSH band gap calculation, assigning a single value to the dielectric screening () for any layered perovskite is particularly challenging as the materials are composed of layers of moieties (i.e., organic and inorganic) with different dielectric constants, hence are not homogeneous. The first, most trivial option is to use the average of the main components of the dielectric tensor (i.e., the average of the diagonal) to define the Hartree exact exchange fraction, that is with . The calculated band gaps within the one-shot DSH() and self-consistent DSH() method for layered perovskites are shown in Figure 3. Similar to 3D CsPbX3, the band gaps calculated employing the PBE and HSE functionals are largely underestimated compared to experiment. The band gap increases from PBE to HSE, PBE0, DSH() and finally DSH(). However, we find that both DSH() and DSH() functionals overestimate the band gaps across all the layered materials irrespective of the crystal symmetry, organic cation, or the type of halogen in the perovskite layer. Taking for example BA2PbBr4 (room temperature phase), DSH0() predicts a band gap of 3.65 eV, which is 0.31 eV higher than the experimental value of 3.34 eV reported by Silver et al. 55 The self-consistent DSH method, DSH(), yields an even higher value of 3.79 eV, exceeding the experimental result by 0.45 eV. This trend is consistent across all layered materials studied here and can be explained as the screening for the DSH0 and DSH is based on , which includes contributions from the inorganic perovskite and the organic layer. In other words, the dielectric screening is an average of a low dielectric constant of the organic layer and a higher one of the inorganic layer. However, for the layered halide perovskites of interest here, the electronic band edges that define the electronic band gap are composed of states of the inorganic layer, as the valence band is mainly Pb 6-orbitals and the halogen atoms -orbitals, while the conduction band minimum is mostly of Pb 6-orbitals and the halogens orbitals 56, 57. Therefore, one can argue that using DSH with the dielectric constant solely of the inorganic perovskite layer is more reasonable for predicting the band gaps accurately.
@@
Considering the cubic phases of 3D halide perovskites allows defining unambiguously . Yet, for layered halide perovskites there are various methods developed to extract the dielectric constants of the inorganic () and organic () layers. For example, can be approximated as the bulk dielectric constant of the respective halide salt PbI2 58. One can employ a dielectric profile calculated using different flavors of DFT to assess and 59, 60, or model the structure using a two capacitor model and fit the dielectric constants along one 54 or both directions 61. While these methods can provide reasonable estimates of the dielectric constants, the challenge remains in accurately defining the model’s parameters, for example, the thickness of the inorganic layers or defining a clean interface between the two layers. Moreover, the accuracy of the dielectric constant depends on the choice of the DFT functional; the semi-local PBE functional, for example, tends to overestimate dielectric constants and gives divergent values for materials with narrow band gaps 36. To overcome this issue in a simple way that is easy to implement in an ab initio parameter-free workflow, we propose employing the DSH using the dielectric constant of the 3D polymorphous perovskites to set the screening parameter. This method corresponds to the band gaps denoted as DSH() shown in Figure 3 and summarized in Table S3. The band gap values predicted within DSH() is a significant improvement with respect to the other functionals. For example, the band gap of BA2PbBr4 using DSH() is 3.26 eV, in excellent agreement with the experimental value of 3.34 eV 55. Interestingly, PBE0 also gives band gaps close to the measured electronic band gaps and similar to DSH(). For the iodides in particular, the similar performance of PBE0 and DSH() can be attributed to a similar Hartree fraction for the long-range screening, which is in PBE0, and in DSH() as is 4.52, and the effects from the short-range screening is less important as the case of 3D halide perovskites shown in Figure 2. Similarly, for bromides, DSH() is slightly above the measured band gaps, and PBE0 slightly below, which can also be attributed to the different screening at the long-range as , yields of 0.28. In contrast, for chlorides an gives an of 0.31. The exact shape of the screening for all layered perovskites studied here is shown in Figure S2.

### v2

@@ Abstract
Accurate prediction of electronic band gaps in halide perovskites using density functional theory is crucial for optoelectronic applications. Standard hybrid functionals like HSE and PBE0 are becoming computationally accessible, yet can fail at predicting the band gaps for three-dimensional (3D) and/or layered halide perovskite. This study evaluates the doubly screened dielectric-dependent hybrid (DSH) functional for predicting band gaps of Pb- and Sn-based inorganic and hybrid 3D halide perovskites, as well as layered hybrid perovskites. The DSH functional employs material-dependent mixing parameters derived from macroscopic dielectric constants, and accurately predicts band gaps for 3D perovskites only if structural local disorder is taken into account. For layered hybrid perovskites, the DSH functional based on average dielectric constants tends to overestimate the band gaps. To improve predictions, we propose using the calculated dielectric constant of the respective 3D perovskites to define the DSH screening. This method is then applied to Pb- and Sn-based layered halide perovskites with various organic spacers and multilayered structures, such as BA2MAn-1PbnI3n+1 with n =1, 2, 3, resulting in improved precision. The HSE functional systematically underestimates band gaps in layered perovskites due to the missing non-local long-range dielectric screening. On the other hand, the PBE0 is in good agreement with the experimental values, in particular for the layered iodide perovskites. The computational framework introduced here provides an efficient parameter-free ab initio methodology suitable for predicting the electronic properties of 3D, layered halide perovskites and their heterostructures, towards modelling materials for advanced optoelectronic devices.
@@
Lead-based hybrid halide perovskites have emerged as versatile materials with tunable structural, electronic, and optical properties due to their unique inorganic corner-sharing octahedral lattices and complex organic components. These diverse material compositions enable a spectrum of functionalities that have been successfully applied in various technological fields, such as solar cells, photodetectors, and light-emitting devices [41, 33, 27, 21]. Halide perovskites are generally classified into three-dimensional (3D) and layered halide perovskites. The 3D halide perovskites are represented by the general formula AMX3, where M is typically a divalent cation in the center of the octahedron such as Pb2+ or Sn2+, X is a halogen ( I, Br, or Cl), and the A site is typically occupied by Cs+ or organic cations such as methylammonium (MA) or formamidinium (FA). In recent years, solar cells based on 3D halide perovskites have achieved remarkable power conversion efficiencies today comparable to those achieved by silicon cells [41]. However, these 3D materials can suffer from poor environmental stability when exposed to air and/or illumination [2]. Their layered counterparts, also called two-dimensional (2D) halide perovskites, are further classified into Ruddlesden-Popper (RP), Dion-Jacobson (DJ), and alternating cation (ACI) structures, depending on the choice of organic molecules that separate the layers and the resulting stacking arrangement [7]. For example, the RP structure typically has two monovalent organic cations and the general formula , where is the number of connected MX6 octahedra in the structure [38]. Layered halide perovskites can exhibit significantly higher material stability than 3D perovskites. However, their photovoltaic performance is usually lower [31, 11]. Within these layered materials, the organic moieties are usually insulating; thus, the charge carriers are primarily constrained to the inorganic layers. This leads to the natural formation of quantum wells (QW), in which the optical band gap increases with decreasing due to quantum and dielectric confinement [6, 18]. Most recently, the combination of the two in the so-called 3D/2D heterostructures has achieved superior stability compared to 3D and higher efficiency than layered halide perovskites, making them a promising platform for optimizing device performance in various architectures [37]. Consequently, knowledge about the energy levels is critical to monitor the band alignment between layered and 3D structures [37].
@@
Density Functional Theory (DFT) calculations, employing semi-local exchange-correlation (XC) functionals like the generalized gradient approximation (GGA) with spin-orbit coupling effects, lead to a sizable and systematic underestimation of the electronic band gap. The state-of-the-art method to go beyond DFT is Hedin’s many-body quasiparticle perturbation theory [24] that accounts for exchange-correlation effects via the self-energy, defined based on the Green’s function, , and the dynamical Coulomb screening potential . The quasiparticle conduction band minimum (CBM) and valence band maximum (VBM) can be interpreted as the ionization energy (IE) measured by ultraviolet photoemission spectroscopy (UPS) and the electron affinity (EA) obtained from inverse photoemission spectroscopy (IPES), respectively. However, the approximation is computationally expensive [57], which limits its use in large systems containing hundreds or thousands of electrons, such as layered halide perovskites and very few works have reported the GW band gaps in the entire structure (i.e., including the organic moieties) of layered halide perovskites [61, 19, 39, 16, 30].
@@
An alternative approach to is using the so-called hybrid functionals within DFT constructed by admixing a non-local Hartree fraction to the local exchange potential [36]. This mixing mitigates the self-interaction error, and has been shown for several semiconductors to result in improved band gaps and reasonable energy level alignment at a computational cost that can be significantly smaller than calculations [10]. Well-known hybrid functionals include PBE0 [42], which adds 25% (i.e., the mixing parameter is set to 0.25) exact Hartree-Fock exchange to the exchange-correlation potential. HSE [25, 29] is similar to PBE0 but introduces a range-separation parameter that partition the Coulomb interaction into short- and long-range components. These hybrid functionals significantly improve standard DFT, yet PBE0 tends to overestimate band gaps, particularly for narrow band gap materials. At the same time, HSE underestimates band gaps of moderate band gap semiconductors and wide band insulators [51]. To overcome this limitation, the mixing parameters and are often modified to reproduce experimental results, which limits the predictive capability of hybrid functionals [23, 14, 55]. A good example of this limitation is the case of halide perovskites, for which hybrid functionals are being employed with a wide range of mixing parameters, in many cases surpassing the 60% of exact exchange [14, 4, 55, 32]. Such manual tuning of the parameters in hybrid functionals is impractical and ineffective if the band gap of the material is not known a priori. It can also be challenging to predict suitable mixing parameters in complex heterostructures. Hence, it is desirable to develop mixing parameters independent of empirical tuning and have an intrinsic physical meaning.
@@
To this end, it has been shown that the inverse macroscopic dielectric constant , which relates to the screening strength in materials, can be employed in a ”parameter-free” dielectric-dependent hybrid functional (DDH) which gives good agreement with experiments for representative materials with a wide range of band gaps and dielectric constants [1, 35, 40]. A fully self-consistent version sc-DDH, which involves evaluating at the end of each DDH cycle, was also developed [49]. A range-separated version (sc-RS-DDH) [50], with the long-range screening set at the empirical value of 0.25, has been proposed that further improves the band gap prediction with respect to sc-DDH. In the same spirit, but without any empirical parameter, two types of dielectric-dependent hybrid functionals have recently been introduced. First, the doubly screened dielectric-dependent hybrid (DSH) functional [12], which accounts for both electronic and metallic screening, and the dielectric-dependent range-separated hybrid functional using the Coulomb attenuated method (DD-RS-CAM) [9]. Although both functionals use the same model dielectric function [8, 46], in each case the range separation parameter is determined slightly differently. In DSH, is derived from modified Thomas-Fermi screening, whereas in DD-RS-CAM is obtained by fitting the first-principles RPA dielectric function with the model dielectric function. Nevertheless, these two functionals are practically the same and do not give rise to significant differences in the predicted band gaps of materials [50, 60]. The model macroscopic dielectric function in DSH and DD-RS-CAM disregards both local field effects (i.e., off-diagonal elements) and dynamical effects and can be viewed as a static approximation to the diagonal Coulomb hole and screened exchange (COHSEX) [24, 20] approximation to the GW self-energy. Overall, the validity of this approximation depends on how well the semi-local correlation represents the Coulomb hole and the significance of the dynamical effects. [9] Table S1 summarizes the parameters of the different hybrid functionals.
@@
To date, DSH has been applied to 3D halide perovskites [5, 59, 60, 58], and, to the best of our knowledge, no reports are validating DSH for the prediction of the band gap of layered halide perovskites, which remains a challenge. Nevertheless, the computational studies on 3D perovskites employ high-symmetry structures and disregard effects arising from structural local disorder, which are critical for the description of halide perovskites, playing an important role on their structural, electronic, optical, and phonon properties [64, 62, 63]. In this work, we compute the electronic band gaps of the most prominent 3D and layered halide perovskites using the DSH hybrid functional. We do so to explore the possibility of developing a consistent workflow that would permit calculating the band gaps of all these materials without having to tune any parameter of the hybrid functionals empirically. To this end, we consider a set of layered halide perovskites, known to pose a challenge for hybrid functionals, containing different organic cation spacers and different halogen atoms (i.e., Cl, Br, and I), as well as a layered perovskite series with n = 2 and n=3. All materials studied in this work are shown in Figure 1. We assess the performance of the DSH functional in both its one-shot (DSH0) and self-consistent (sc-DSH) forms, comparing the results with those obtained using typical hybrid functionals, such as HSE and PBE0, and experimental measurements. We also explore the importance of the hybrid functionals’ short- and long-range separation on band gap for each system.
@@
We begin by analyzing how the doubly screened hybrid (DSH) functional performs on the prediction of the band gaps of prototypical cubic structures of 3D halide perovskites AMX3 (A = Cs, MA, FA; M = Sn, Pb; X = Cl, Br, I). We model the structures using the lattice parameters that are summarized in Table S2. All computed band gaps include spin-orbit coupling effects, important in halide perovskites [15]. The structures of the organic-inorganic halide perovskites MAPbCl3 and FAPbCl3 are taken from Ref. 43 and were generated using the ASDM method. [62, 63]. We compare these predictions with those obtained using PBE, PBE0, and HSE06, as well as the measured experimental electronic band gaps (See Table S2). Figure 2 shows the band gaps calculated employing different functionals (bars) and the measured band gaps (solid lines). As an additional reference, we include the standard semiconductors Si and GaAs, which have similar band gaps yet dielectric constants much higher than AMX3. Not surprisingly, all hybrid functionals drastically improve the PBE band gap values. For the standard semiconductors, HSE values are very close to the experimental band gaps, and so are the values obtained from the dielectric-dependent DSH0 and DSH. On the other hand, PBE0 significantly overestimates the band gap as range separation is critical for the functional form in both Si and GaAs. Moving on to the cubic phases of the 3D halide perovskites AMX3, the hybrid functionals fail to describe the band gaps as all functionals significantly underestimate the measured band gaps. However, this underestimation is related to neglecting local disorder (also referred to as positional ”polymorphism”), which is critical for the accurate description of the electronic structure of halide perovskites at higher temperatures, including their dielectric properties [65, 64, 62, 63]. Some of us have previously shown that such local disorder (shown in the inset structures of Figure 2) has significant effects on the band gaps, lattice dynamics, and electron-phonon coupling of halide perovskites [62, 63]. Therefore, we consider the locally disordered configurations p-AMX3 to assess the performance of the hybrid functionals. As shown in the last panel of Figure 2, we find a significant increase of the band gap values in p-CsPbX3 resulting from the local disorder for all the functionals, ranging from 570 meV (PBE, p-CsPbCl3) to 960 meV (DSH0, p-CsPbI3), leading to a better agreement with the measurements.
@@
Next, we discuss the effects of self-consistency on improving DSH band gaps. In both full-range [49] and range-separated dielectric dependent hybrid functionals [12, 9], early results indicated that self-consistency improves the band gaps values compared to one-shot methods, especially for narrow band gap materials [12]. However, such self-consistency worsens the results for transition metal oxides, which is attributed to the underestimated dielectric constants [34]. Here, we observe that self-consistency leads to relatively minor changes in the calculated band gaps, with the change being marginal in the case of Si, GaAs, and p-CsPbI3. Figure S3 shows the band gap variation and dielectric constants for each DSH iteration. We also note that the method employed to evaluate in the self-consistency cycle can be critical. can be computed using different methods, including density functional perturbation theory (DFPT) [3], finite electric field in the Berry phase approximation [44], or linear optics in the Random Phase Approximation (RPA) [17], each with different advantages and limitations discussed in the supporting information file. Here, we employ the finite electric field method, as detailed in the computational details, to include local field effects, which are important for halide perovskites. However, converging remains challenging when employing hybrid functionals.
@@
Next, we employ the DSH functional to set up a parameter-free workflow for layered halide perovskite materials. To this end, we consider a series of prototypical lead-based layered structures with various halogen atoms (i.e., Cl, Br, I) and different organic spacers, as seen in Table S3. Figure 1d-g shows the materials with the organic cations BA (butylammonium), PMA (phenylmethyl ammonium), PEA (phenylethylammonium) and FPEA (4F-phenylethylammonium), respectively. We have not considered locally disordered structures for the layered perovskites, as our calculations are based on experimental low-symmetry structures with significant distortion and therefore, the effects are expected to be small. For example, the change in total energy due to structural local disorder for orthorhombic CsPbI3 is reported at just 1 meV per formula unit [62]. Also, in the absence of experimental electronic band gaps, we correct the experimental optical band gaps (E) using the empirical relation of the exciton binding energy (), derived for layered perovskites (i.e., ). [22] The ab initio computation of the exciton binding energy would require to solve the GW plus Bethe-Salpeter equation (GW+BSE). To date, only a handful of studies have used the GW+BSE method to compute band gap and exciton binding energy in the full structure (i.e., including organic moieties) of layered halide perovskites [61, 19, 39, 16, 30]. For example, Yin et al. [61] investigated the origin of light emission in two layered perovskites. Giorgi et al. [19] calculated the optoelectronic properties of isolated and layered with n=1 and n=2. Filip et al. [16] analyzed the effects of various organic cations on the band gaps, dielectric screening, and exciton binding energies of a set of A2PbBr4 compounds.
@@
For the DSH band gap calculation, assigning a single value to the dielectric screening () for any layered perovskite is particularly challenging as the materials are composed of layers of moieties (i.e., organic and inorganic) with different dielectric constants, hence are not homogeneous. The first, most trivial option is to use the average of the main components of the dielectric tensor (i.e., the average of the diagonal) to define the Hartree exact exchange fraction, that is with . The calculated band gaps within the one-shot DSH() and self-consistent DSH() method for layered perovskites are shown in Figure 3. Similar to 3D CsPbX3, the band gaps calculated employing the PBE and HSE functionals are largely underestimated compared to experiment. The band gap increases from PBE to HSE, PBE0, DSH() and finally DSH(). However, we find that both DSH() and DSH() functionals overestimate the band gaps across all the layered materials irrespective of the crystal symmetry, organic cation, or the type of halogen in the perovskite layer. Taking for example BA2PbBr4 (room temperature phase), DSH0() predicts a band gap of 3.65 eV, which is 0.31 eV higher than the experimental value of 3.34 eV reported by Silver et al. [47] The self-consistent DSH method, DSH(), yields an even higher value of 3.79 eV, exceeding the experimental result by 0.45 eV. This trend is consistent across all layered materials studied here and can be explained as the screening for the DSH0 and DSH is based on , which includes contributions from the inorganic perovskite and the organic layer. In other words, the dielectric screening is an average of a low dielectric constant of the organic layer and a higher one of the inorganic layer. However, for the layered halide perovskites of interest here, the electronic band edges that define the electronic band gap are composed of states of the inorganic layer, as the valence band is mainly Pb 6-orbitals and the halogen atoms -orbitals, while the conduction band minimum is mostly of Pb 6-orbitals and the halogens orbitals [13, 54]. Therefore, one can argue that using DSH with the dielectric constant solely of the inorganic perovskite layer is more reasonable for predicting the band gaps accurately. Figures S4 and S5 show the DSH() density of states (DOS) for all layered perovskites presented in Figure 3 and Figure 4a, respectively, showing that the band edges are primarily composed of orbitals from the inorganic framework.
@@
Considering the cubic phases of 3D halide perovskites allows defining unambiguously . Yet, for layered halide perovskites there are various methods developed to extract the dielectric constants of the inorganic () and organic () layers. For example, can be approximated as the bulk dielectric constant of the respective halide salt PbI2 [26]. One can employ a dielectric profile calculated using different flavors of DFT to assess and [45, 56], or model the structure using a two capacitor model and fit the dielectric constants along one [22] or both directions [48]. While these methods can provide reasonable estimates of the dielectric constants, the challenge remains in accurately defining the model’s parameters, for example, the thickness of the inorganic layers or defining a clean interface between the two layers. Moreover, the accuracy of the dielectric constant depends on the choice of the DFT functional; the semi-local PBE functional, for example, tends to overestimate dielectric constants and gives divergent values for materials with narrow band gaps [12]. To overcome this issue in a simple way that is easy to implement in an ab initio parameter-free workflow, we propose employing the DSH using the dielectric constant of the 3D polymorphous perovskites to set the screening parameter. This method corresponds to the band gaps denoted as DSH() shown in Figure 3 and summarized in Table S3. The band gap values predicted within DSH() is a significant improvement with respect to the other functionals. For example, the band gap of BA2PbBr4 using DSH() is 3.26 eV, in excellent agreement with the experimental value of 3.34 eV [47]. Interestingly, PBE0 also gives band gaps close to the measured electronic band gaps and similar to DSH(). For the iodides in particular, the similar performance of PBE0 and DSH() can be attributed to a similar Hartree fraction for the long-range screening, which is in PBE0, and in DSH() as is 4.52, and the effects from the short-range screening is less important as the case of 3D halide perovskites shown in Figure 2. Similarly, for bromides, DSH() is slightly above the measured band gaps, and PBE0 slightly below, which can also be attributed to the different screening at the long-range as , yields of 0.28. In contrast, for chlorides an gives an of 0.31. The exact shape of the screening for all layered perovskites studied here is shown in Figure S2.
In order to quantify the performance of the different functionals in describing the band gap of layered halide perovskites, we calculate the mean absolute error (MAE) and root mean square error (RMSE). Figure 4a summarizes the performance of all hybrid functionals for all the layered materials, as well as by halogen (A2PbCl4, A2PbBr4, A2MI4 (M=Sn,Pb)). The HSE functional is the least suitable, with errors close to 1.0 eV. PBE0, on the other hand, predicts the band gap reasonably well across the layered materials, particularly for iodides.